OpenStudy (anonymous):

write the solution using set builder notation. 8 - 9 (2x +4) <= -2x + 13

OpenStudy (anonymous):

http://www.virtualnerd.com/algebra-1/algebra-foundations/inequality-in-set-builder-notation.php \[8-18x-36 \le -2x+13\] \[-28 \le 16x+13\] \[-41 \le 16x\] \[x \ge \frac{-41}{16}\] {x|x>=-41/16}